$SPAD/src/input isprime.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool isprime.output
)set message test on
)set message auto off
)clear all

———-

The first two (2152302898747 and 3474749660383) are the only strong pseudoprimes bases 2,3,5,7,11 less than 10^{1}3: the remaining three are the only Carmichael numbers less than 10^{1}6 with this property.

— * —

--S 1 of 15
n := 6763*10627*29947
--R
--R
--R   (1)  2152302898747
--R                                                        Type: PositiveInteger
--E 1
--S 2 of 15
prime?(n)
--R
--R
--R   (2)  false
--R                                                                Type: Boolean
--E 2
--S 3 of 15
factor(n)
--R
--R
--R   (3)  6763 10627 29947
--R                                                      Type: Factored(Integer)
--E 3
--S 4 of 15
n := 1303*16927*157543
--R

--R
--R   (4)  3474749660383
--R                                                        Type: PositiveInteger
--E 4
--S 5 of 15
prime?(n)
--R
--R
--R   (5)  false
--R                                                                Type: Boolean
--E 5
--S 6 of 15
factor(n)
--R
--R
--R   (6)  1303 16927 157543
--R                                                      Type: Factored(Integer)
--E 6
--S 7 of 15
n := 3739*18691*153259
--R
--R
--R   (7)  10710604680091
--R                                                        Type: PositiveInteger
--E 7
--S 8 of 15
prime?(n)
--R
--R
--R   (8)  false
--R                                                                Type: Boolean
--E 8
--S 9 of 15
factor(n)
--R
--R
--R   (9)  3739 18691 153259
--R                                                      Type: Factored(Integer)
--E 9
--S 10 of 15
n := 46411*232051*417691
--R
--R
--R   (10)  4498414682539051

--R                                                        Type: PositiveInteger
--E 10
--S 11 of 15
prime?(n)
--R
--R
--R   (11)  false
--R                                                                Type: Boolean
--E 11
--S 12 of 15
factor(n)
--R
--R
--R   (12)  46411 232051 417691
--R                                                      Type: Factored(Integer)
--E 12
--S 13 of 15
n := 21319*106591*3005839
--R
--R
--R   (13)  6830509209595831
--R                                                        Type: PositiveInteger
--E 13
--S 14 of 15
prime?(n)
--R
--R
--R   (14)  false
--R                                                                Type: Boolean
--E 14
--S 15 of 15
factor(n)
--R
--R
--R   (15)  21319 106591 3005839
--R                                                      Type: Factored(Integer)
--E 15
)spool
)lisp (bye)

———-

References^{[1] nothing}

6